OpenStudy (demaje5):

Alan found the value of the expression (15 - 6)2 to be 189. Is he right?

OpenStudy (shadelily):

Everything that is in brackets has to be done first so 15 - 6 = 9 Then do the rest. 9 * 2 = 18 not 189, thus the answer is that Alan is not right.

OpenStudy (asc.bchs):

Another way to do this is to distribute the two. 2(15 - 6) = 30 - 12 = 18. Same answer that @ShadeLily got, but different way to do it. Either way, Alan is wrong.

OpenStudy (mathstudent55):

My guess is that the 2 is an exponent, but of course, I can't be sure. \(\Large (15 - 6)^2\) Correct: \(\Large (15 - 6)^2 = 9^2 = 81\) Incorrect (maybe what Alan did): \(\Large (15 - 6)^2 \color{red}{=} 15^2 - 6^2 = 225 - 36 = 189\) The first equal sign above is in red because it is incorrect. That is the step that Alan did wrong.

OpenStudy (mathmale):

In the future, please use the symbol " ^ " to indicate exponentiation. (15-6)2 is ambiguous. You want (15-6)^2.

mathmale:

Agreed. And then 9^2=?